OpenStudy (anonymous):

What is the value of the y-coordinate of the solution to the system of equations 2x=y=8 and x-3y=-3

OpenStudy (anonymous):

Is that supposed to be 2x-y=8 ?

OpenStudy (anonymous):

yes my mistake sorry

OpenStudy (anonymous):

2x-y=8 equation I x-3y=-3 equation II Multiply equation II by 2 2x-y=8 equation III 2x-6y=-6 equation IV Subtract equation IV from equation III to get rid of the 2x -y-(-6y)=8-(-6) -y+6y=8+6 5y=14 y=15/4

OpenStudy (anonymous):

Hope that helps

OpenStudy (anonymous):

Yes it does thank you

OpenStudy (anonymous):

No problem :)